OpenStudy (anonymous):

Select the equation of a line that is parallel to the line on the graph and passes through the point (2, 3).

OpenStudy (mathmale):

Note: two lines that are parallel have the same slope. So please go to the illustration, look at the line given there, and determine its slope. Then use the point-slope formula: \[y-y _{1}=m(x-x _{1})\] and simplify the result.

OpenStudy (anonymous):

still not understanding @mathmale

OpenStudy (mathmale):

Andrew: I'd bet there was an illustration that came with this problem. Right? if so, please share that illustration. We have to obtain the slope of "the given line" from that illustration.

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Thank you. I see that this line has a negative slope. Please determine the slope, m, of this line.

OpenStudy (anonymous):

-2

OpenStudy (mathmale):

You are definitely on the right track. Note, however, that the slope is\[\frac{ \Delta y }{ \Delta x }\], and you've put in delta y and delta x upside down. Try again, please. you're getting there!

OpenStudy (anonymous):

-1/-2 is this right

OpenStudy (mathmale):

You can see that the graph slopes downward, so the slope is negative. Thus, -1/(-2) could not be right. Instead, m = ??

OpenStudy (anonymous):

2

OpenStudy (mathmale):

Andrew: The "rise" is actually negative: It's -1. The "run" is positive: It's 2. therefore the slope of this line is m = ??

OpenStudy (anonymous):

- one half

OpenStudy (anonymous):

if this is wrong im really sorry for wasting your time but I really don't get this

OpenStudy (mathmale):

that's correct. Thanks for y our patience. your m is then m=-1/2. this new line passes thru the point (2,3). Use the point-slope form of the equation of a straight line to find the equation of this new line.